A brace for a shelf has the shape of a right triangle. Its hypotenuse is 10 inches long and the two legs are equal in length. How long are the legs of the triangle?
Answer by oberobic(2304)

Use the Pythagorean Theorem.
c^2 = 10^2 = 100
c^2 = a^2 + b^2
a = b
100 = a^2 + a^2 = 2*a^2
a^2 = 50
a = sqrt(50) = sqrt(25*2) = 5*sqrt(2)
so,
The legs are each 5*sqrt(2) inches.